OpenStudy (anonymous):

Solve 2/5x = -3/15 + 1/x

OpenStudy (anonymous):

\[\frac{2 x}{5}=-\frac{1}{5}+\frac{1}{x}\]Move the RHS to the LHS and combine fractions.\[\frac{-5+x+2 x^2}{5 x}=0\]Solve the numberator set to zero for x.\[\left\{x=\frac{1}{4} \left(-1-\sqrt{41}\right),x=\frac{1}{4} \left(-1+\sqrt{41}\right)\right\} \]

OpenStudy (anonymous):

numberator should have been spelled, numerator.